# Computing the Matching Distance of $2$-Parameter Persistence Modules from Critical Values

### v2

@@ 4 Existence of the switch points
Recall the notation $p_{L}(u)$ for the parameter value of the push of a point $u$ onto the line $L$, and recall that, for every $u\in\mathbb{R}^{2}$, $u$ pushes rightwards if $A^{L}_{u}=\{2\}$, upwards if $A^{L}_{u}=\{1\}$, and to itself if $A^{L}_{u}=\{1,2\}$. In Lemma 4.2 we will also use the following notation used to indicate sets of lines which partition a particular configuration of points in a particular way.
@@
  • $\beta=0$ and at least one of $\alpha_{1},\alpha_{2}$ equal 00,
@@
$\Delta_{L}(u,v,w,x;\delta,\eta)>0\textrm{ and }\Delta_{L^{\prime}}(u,v,w,x;\delta,\eta)<0.$
@@ 5 The matching distance algorithm
Because this has to be repeated for each line, whose number is $O(r^{8})$, we get the total time complexity for step 4 bounded by $O(r^{8+\mu})$, which dominates that of the previous steps.
In conclusion, the time complexity of our method is roughly $O(r^{11})$, which agrees with that of [25] and is higher than that of [5] which is $O(r^{5}\log^{3}r)$.
Compared to the algorithms of [20, 5], the above steps avoid applying the point-line duality, thus giving a more straightforward intuition on the lines that attain the matching distance. For implementation purposes, we note that the matching distance computation is parallelizable, with bottleneck distance computations carried out in parallel on different lines. Additionally, the computation of candidate switch points is also parallelizable based on the cases described in the proof of Lemma 4.2, item 2 through item 4. Algorithms to prune the set of candidate switch points are detailed in [8].
6 Conclusion
@@
  • Theoretical methods for computing the matching distance between two $2$-parameter persistence modules, which can be used to produce algorithms with comparable time complexity to [20];
@@
We expand on these bullets in order. First, one advantage of the theoretical framework developed in this paper is that it allows us to formulate explicit ways of computing the switch-points from the critical values of two $2$-parameter persistence modules (see [8]). This allows us to provide an algorithm performing the exact computation of the matching distance in two dimensions as the maximum over a finite set of lines, as guaranteed by Theorem 3.1.
@@
About diagonal lines, we show via Example 2.11 that in some cases, lines through pairs of critical points are not sufficient in order to compute the matching distance, but adding lines through critical points with the diagonal direction may turn out to be sufficient. This agrees with the observation in [13] that diagonal lines play a key role in the study of $2$-parameter persistence. However, Examples 2.12 and 2.13 show that adding in diagonal lines is also not always sufficient as lines through critical and switch points need to be considered. While critical points are always proper points, our theoretical results show that some of switch points are points at infinity of the projective plane. This way it is possible to limit the number of slopes of lines on which the matching distance can be obtained.
About vertical and horizontal lines, we are able to give an explicit formula for the computation of the weighted bottleneck distance along horizontal and vertical lines. Moreover, we are able to conclude that these lines can be excluded from the computation of the overall matching distance between two $2$-parameter persistence modules — thanks to the inclusion of diagonal lines in the computation, there is always a line with positive but finite slope for which the weighted bottleneck distance is greater than that along a horizontal or vertical line. Viewing the parameter space as a subset of the projective plane helps in understanding the importance of vertical, diagonal, and horizontal lines, and makes the treatment of these lines more consistent.
Third, in [20, 5], the authors leverage the duality between points and lines in the plane to compute the matching distance, performing computations in the dual plane. In comparison, our method is directly defined in the parameter space of the persistence modules, helping the understanding of the finite set of lines involved in the maximization for the computation of the matching distance. In particular, identifying the parameter space with a subset of the projective plane helps to highlight the importance of both proper and infinite switch points (switch points which represent positive slopes).
@@ Acknowledgements
E.S. is supported by European Research Council (ERC) grant no. 788183, ‘Alpha Shape Theory Extended’.
@@
We also thank Martin Hafskjold Thoresen for help with computation.

### v3

@@ 4 Existence of the switch points
Recall the notation $p_{L}(u)$ for the parameter value of the push of a point $u$ onto the line $L$, and recall that, for every $\in\mathbb{R}^{2}, uu$ pushes rightwards if $A^{L}_{u}=\{2\}$, upwards if $A^{L}_{u}=\{1\}$, and to itself if $A^{L}_{u}=\{1,2\}$. In Lemma 4.2 we will also use the following notation used to indicate sets of lines which partition a particular configuration of points in a particular way.
@@
  • $\beta=0$ and at least one of $\alpha_{1},\alpha_{2}$ equal 0,
@@
[…]
@@
4.1 Connections to related works
The need to consider switch points is ubiquitous in the literature about the computation of the matching distance. To the best of our knowledge, however, switch points’ coordinates are always left implicit and never worked out. See, for example, [18, Definition 5.1] and equations (1) in [25].
Working out the explicit formulas, as done in the proof of Lemma 4.2, we found it convenient to pass to the projective plane because some switch points belong to the line at infinity. In other words, a switching may happen when lines pass through a given direction rather than a given point. In the dual plane, this corresponds to a point passing through a vertical line. To the best of our understanding, this case is not singled out in the arrangement $\mathcal{A}$ of [25]. The only vertical line explicitly discussed in the dual plane is the one corresponding to the diagonal direction in the primal plane.
Interestingly, vertical lines in the dual plane play a role in [7]: The change through the value $\lambda$ of the weighted bottleneck distance is shown to happen at a dual point on a line in dual space that can be vertical and therefore does not have a corresponding point in primal space (if not at infinity).
@@ 5 The matching distance algorithm
Because this has to be repeated for each line, whose number is $O(r^{8})$, we get the total time complexity for step 4 bounded by $O(r^{8+\mu})$, which dominates that of the previous steps. In conclusion, the time complexity of our method is roughly $O(r^{11})$.
5.1 Connections to related works
The algorithm we streamlined above mirrors, in the primal plane, the algorithm of [25]. The differences are that we manage not to consider horizontal lines, and that we have explicit formulas for switch points. Computationally, the running time agrees with that of [30] and is higher than that of [7] which is $O(r^{5}\log^{3}r)$.
We note that the matching distance computation is parallelizable, with bottleneck distance computations carried out in parallel on different lines. Additionally, the computation of candidate switch points is also parallelizable based on the cases described in the proof of Lemma 4.2, item 2 through item 4. Algorithms to prune the set of candidate switch points are detailed in [10].
6 Conclusions
@@
  • Theoretical results for computing the matching distance between two $2$-parameter persistence modules, in a combinatorial setting and while staying in the primal plane;
@@
We expand on these bullets in order. First, one advantage of the theoretical framework developed in this paper is that it allows us to formulate explicit ways of computing the switch-points from the critical values of two $2$-parameter persistence modules (see [10]). This allows us to prove that the matching distance in two dimensions as the maximum over a finite set of lines, as guaranteed by Theorem 3.1.
@@
About diagonal lines, we show via Example 2.12 that in some cases, lines through pairs of critical points are not sufficient in order to compute the matching distance, but adding lines through critical points with the diagonal direction may turn out to be sufficient. This agrees with the observation in [15] that diagonal lines play a key role in the study of $2$-parameter persistence. However, Examples 2.13 and 2.14 show that adding in diagonal lines is also not always sufficient as lines through critical and switch points need to be considered. While critical points are always proper points, our theoretical results show that some switch points are points at infinity of the projective plane. This way, it is possible to limit the number of slopes of lines on which the matching distance can be obtained.
Regarding vertical and horizontal lines, we give an explicit formula for the computation of the weighted bottleneck distance along horizontal and vertical lines. Moreover, we conclude that these lines can be excluded from the computation of the overall matching distance between two $2$-parameter persistence modules — thanks to the inclusion of diagonal lines in the computation, there is always a line with positive but finite slope for which the weighted bottleneck distance is greater than that along a horizontal or vertical line. Viewing the parameter space as a subset of the projective plane helps in understanding the importance of vertical, diagonal, and horizontal lines and makes the treatment of these lines more consistent.
Third, in [25, 7], the authors leverage the duality between points and lines in the plane to compute the matching distance, performing computations in the dual plane. In comparison, our method is directly defined in the parameter space of the persistence modules, helping the understanding of the finite set of lines involved in the maximization for the computation of the matching distance. In particular, identifying the parameter space with a subset of the projective plane helps to highlight the importance of both proper and infinite switch points (switch points that represent positive slopes).
@@
So, In any case, it remains as an open question whether the matching distance for $n$-parameter persistence modules with $n\geq 3$ is achieved as a maximum.
@@ Acknowledgements
E.S. was supported for travel to BIRS by European Research Council (ERC) grant no. 788183, ‘Alpha Shape Theory Extended’.
@@
We also thank Martin Hafskjold Thoresen for help with computations, and the anonymous referee who pointed out useful connections to the literature.